OpenStudy (ladiesman217):

MEDAL AND FANS AWARDED help please i dont understand this? The equations only has 2 variables and i thought a quadratic equation requires 3 variables? Use the quadratic formula to solve -6x^2 + 6x = 0

OpenStudy (anonymous):

x= 0 & 1

OpenStudy (anonymous):

it has only one variable, and that variable is x

OpenStudy (ladiesman217):

but how would i set it up in the quadratic formula though

OpenStudy (potatoes.ramu):

x^2=x x^2-x=0 x(x-1)=0 So, x=0 or x=1.

OpenStudy (ladiesman217):

ok thnx

OpenStudy (potatoes.ramu):

no problem :)